app 3 a. Determine the present value of a single $13,100 cash flow in 7 years...

app 3

  1. a. Determine the present value of a single $13,100 cash flow in 7 years if the interest (discount) rate is 8% per year. Round your answer to the nearest cent, if rounding is required.
    $

    b. Determine the number of periods for which $5,820 must be invested at an annual interest (discount) rate of 7% to produce an investment balance of $10,000. Round your answer to the nearest whole number of periods, if rounding is required.
    periods

    c. Determine the size of the annual cash flow for a 25-year annuity with a present value of $49,113 and an annual interest rate of 9%. One payment is made at the end of each year. Round your answer to the nearest cent, if rounding is required.
    $

    d. Determine the annual interest rate at which an investment of $2,542 will provide for a single $4,000 cash flow in 4 years. Round your answer to the nearest whole percentage rate (for example, 10.8% rounds to 11%).
    %

    e. Determine the annual interest rate earned by an annuity that costs $17,119 and provides 15 payments of $2,000 each, one at the end of each of the next 15 years. Round your answer to the nearest whole percentage rate (for example, 10.8% rounds to 11%).
Answer #1

a) =PV(8%,7,,13100) =7643.72

b) =NPER(7%,,-5820,10000) =8 years

c) =PMT(9%,25,49113) =5000.01

d) =RATE(4,,-2542,4000) =12.00%

e) =RATE(15,2000,-17119) =8.00%